# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ 4.2 Temporal Order Regularization Loss
We compute verb/object logits using cosine similarity between visual features and their corresponding text features. We apply standard cross-entropy losses for verb and object prediction: $L_{\text{com}}=L_{V}+L_{O}$.
Composition loss.
Following prior work [16], we combine marginal scores $(\mathbf{p}_{V},\mathbf{p}_{O})$ with conditional logits $(\mathbf{p}_{V|O},\mathbf{p}_{O|V})$ to obtain the composition prediction $\hat{\mathbf{y}}$ as $\hat{\mathbf{y}}=\texttt{Softmax}(\mathbf{p}_{O}\mathbf{p}_{V|O}+\mathbf{p}_{V}\mathbf{p}_{O|V})$. We employ the cross-entropy loss: $L_{\text{comp}}=L_{\text{CE}}(\hat{\mathbf{y}},\mathbf{y})$.
Margin loss: mitigating co-occurrence bias.
Highly frequent verb–object pairs often dominate predictions, even for unseen compositions. To counter this, we introduce a margin loss that penalizes frequent-but-incorrect compositions whose logits approach the ground-truth logit:
$L_{M}=\sum_{c\in\mathcal{C}_{\text{margin}}}\max\left(0,\log\hat{\mathbf{y}}(c)-\log\hat{\mathbf{y}}(g)+m\right),$ (6)
where $g\in\{1,\dots,|\mathbb{Y}^{V}|\times|\mathbb{Y}^{O}|\}$ denotes the ground-truth index, and $\mathcal{C}_{\text{margin}}$ is the intersection of top-$K$ incorrect predictions and the set of frequently co-occurring compositions. This term explicitly discourages reliance on dataset co-occurrence priors.
Total loss.
@@
We train RCORE with the combined objective $L_{\text{total}}$ defined as $L_{\text{total}}=\alpha L_{\text{com}}+\beta L_{\text{comp}}+\gamma L_{\text{TORC}}+\delta L_{M}$ , where $\alpha$, $\beta$, $\gamma$, and $\delta$ control the contribution of each term.
@@ 5 Experimental Results
In this section, we conduct rigorous experiments to address the following research questions: (1) Do conventional ZS-CAR evaluation protocols faithfully reveal true model behavior? If not, what is a more appropriate protocol for exposing real generalization performance? (Section 5.1) (2) Does the proposed compositional gap (CG) metric reveal whether a model actually benefits from compositional understanding beyond independent verb/object predictions? (Section 5.2 and Section 5.3) (3) Does RCORE mitigate the two key failure modes—co-occurrence overfitting and object-driven verb shortcuts—identified in our diagnosis? (Section 5.2) (4) Do these improvements translate to consistent gains across datasets? (Section 5.3) (5) How does each proposed component contribute to robust compositional learning? (Section 5.4) We describe the experimental setup and then report results aligned with these questions.
Table 2: Sth-com [16] results. We show the top-1 verb, object, and composition classification accuracy (%). All models use CLIP-B/16 [28] as the backbone. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Model Bias Calib. Verb Object Composition
Seen Unseen H.M. Seen Unseen H.M. Open-world Closed-world
Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M. H.M. AUC
AIM [37] $\times$ 54.99 39.19 45.76 67.19 53.43 59.53 40.24 (+3.29) 18.50 (-2.44) 25.35 - -
C2C w/o conditioning $\times$ 62.53 48.67 54.74 66.90 52.66 58.93 44.95 (+3.12) 23.94 (-1.69) 25.35 - -
C2C [16] $\times$ 63.60 54.36 58.62 67.72 56.10 61.36 46.31 (+3.24) 30.08 (-0.42) 36.47 - -
RCORE $\times$ 65.93 57.29 61.31 67.82 55.56 61.08 48.56 (+3.85) 32.23 (+0.40) 38.75 - -
AIM [37] 50.32 43.19 46.48 64.02 54.78 59.04 33.40 (+1.19) 24.60 (+0.94) 28.33 34.44 16.99
C2C w/o conditioning 57.08 50.31 53.48 65.07 53.32 58.61 39.32 (+2.18) 26.70 (+1.89) 31.80 37.62 19.74
C2C [16] 60.87 56.81 58.77 65.83 57.37 58.93 42.27 (+2.20) 34.22 (+1.30) 37.82 42.75 23.77
Jung et al. [14] - - - - - - - - - 42.60 23.50
RCORE 63.57 58.77 61.08 66.73 56.78 61.35 45.34 (+2.91) 35.26 (+1.89) 39.67 45.45 26.52
C2C (Baseline)
RCORE (Ours)
Figure 6: RCORE mitigates object-driven shortcuts in verb learning. We visualize confusion matrices for six representative verbs to compare the ability of RCORE and C2C to distinguish opposite temporal semantics on unseen compositions of the Sth-com [16] test set. All values in the confusion matrices are normalized frequencies across the entire verb classes in the dataset.
@@ 5.1 Experimental setup
Existing ZS-CAR works [16, 13, 14] rely on the closed-world protocol, where inference is restricted to a small set of compositions that appear in the validation or test split. This protocol hides a model’s incorrect tendency to over-predict seen compositions and obscures co-occurrence-driven behaviors. A more serious methodological flaw is the use of test-set–tuned bias calibration. Selecting the bias term with access to test labels systematically inflates unseen accuracy and prevents meaningful comparison across methods.
Our evaluation setup.
@@
To expose genuine generalization, we adopt an open-world, unbiased protocol by default: (i) inference spans the full space $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, and (ii) no test labels are used to tune biases. We also report: (a) open-world biased results using a realistic validation set-tuned bias, and (b) closed-world H.M. and AUC for comparison with existing methods. All results in Section 3 also follow the default open-world, unbiased protocol.
@@
Datasets. We evaluate on two ZS-CAR benchmarks: Sth-com [16] and our curated EK100-com. Sth-com is derived from Something-Something V2 [12] and Something-Else [22], containing 79K videos with 161 verbs and 248 objects. Following prior work [39, 31], we also evaluate on the Temporal subset consisting of verbs requiring temporal reasoning. We additionally introduce the EPIC-KITCHENS-100-composition (EK100-com) benchmark, constructed by repurposing EPIC-KITCHENS-100 [7] following the Sth-com protocol. EK100-com includes 71K egocentric videos, 82 verbs, and 228 objects. Full construction details are provided in the Section A.
Evaluation metrics.
We report top-1 composition, verb, and object accuracy on seen and unseen compositions, along with their harmonic mean (H.M.). We also measure the compositional gap $\Delta_{\mathrm{CG}}$ (1) which indicates whether a model benefits from compositional reasoning beyond independent verb/object predictions. (see Section 3) A higher $\Delta_{\mathrm{CG}}$ reflects stronger compositional synergy.
Baselines.
@@
We primarily compare RCORE against C2C [16], the current SOTA ZS-CAR method. To isolate the effect of conditional learning, we also evaluate two independent baselines. AIM [37] is the CLIP [28]-based spatio-temporal adapter model used as our backbone. C2C w/o conditioning augments AIM with all architectural components of C2C except the conditional modeling itself. We additionally include closed-world results of a recent ZS-CAR approach [14].
Table 3: EK100-com results. We show the top-1 verb, object, and composition classification accuracy (%). All models use CLIP-B/16 [28] as the backbone. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Model Bias Calib. Verb Object Open-world Composition
Seen Unseen H.M. Seen Unseen H.M. Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M.
AIM [37] $\times$ 60.04 39.78 47.85 56.04 46.33 50.72 38.60 (+4.95) 14.74 (-3.69) 21.34
C2C [16] $\times$ 66.19 49.71 56.78 56.61 47.48 51.64 42.72 (+5.25) 22.38 (-1.22) 29.38
RCORE $\times$ 66.07 52.07 58.24 57.76 47.11 51.89 43.86 (+5.70) 24.24 (-0.29) 31.23
AIM [37] 52.42 49.54 50.94 50.70 51.05 50.87 28.00 (+1.42) 32.07 (+6.78) 29.89
C2C [16] 62.21 57.74 59.89 51.66 54.16 52.88 36.05 (+3.91) 37.96 (+6.69) 36.98
RCORE 62.47 59.77 61.09 53.42 54.79 54.10 36.53 (+3.16) 39.78 (+7.03) 38.09
@@ 5.2 Analysis
In this section, we validate that our RCORE framework effectively addresses the two key challenges in ZS-CAR—(i) reliance on co-occurrence statistics and (ii) object-driven shortcuts in verb learning—and achieves improved compositional representations compared to the baseline [16].
Overcoming reliance on co-occurrence statistics.
@@
In Figure 5 (a), we present the False Co-occurrence Prediction (FCP) curves of RCORE and the baseline [16] on Sth-com. While the baseline’s FCP ratio increases substantially during training (11% $\rightarrow$ 17%), the FCP of RCORE decreases (9% $\rightarrow$ 8%), indicating that RCORE is markedly less dependent on co-occurrence statistics.
Addressing object-driven shortcuts.
In Figure 5 (b), we compare the cosine similarity between the original verb features $\mathbf{f}^{V}$ and the reversed verb features $\mathbf{f}^{V}_{\text{rev}}$ for RCORE and the baseline [16] over the course of training. While the baseline [16] maintains a high cosine similarity ($0.91$) due to its reliance on object-driven shortcuts, RCORE drives this value to a strongly negative (0.73-0.73), successfully distinguishing opposite temporal semantics such as ’opening’ vs.’closing’.
In Figure 6, we further visualize this behavior through verb confusion matrices on the Sth-com [16] test set. RCORE cleanly separates pairs of opposite verbs, while C2C [16] frequently confuses them, e.g. misclassifying ‘unfolding’ as ‘folding’. These results collectively demonstrate that C2C fails to overcome the critical challenge of object-driven shortcuts in verb learning, whereas RCORE directly targets this failure mode and learns verb representations that generalize robustly to unseen compositions.
The learned verb representations are indeed effective.
@@
We further assess the quality of the learned verb representations by evaluating temporal modeling ability on the Temporal subset of Sth-com [16], as shown in Figure 5 (c). To measure a model’s dependence on static cues, we compare its performance when using temporally shuffled verb features $\mathbf{f}^{V}_{\text{shuffled}}$ versus the original features $\mathbf{f}^{V}$. Across both seen and unseen compositions, RCORE exhibits a substantially larger performance gap between original and shuffled features. This demonstrates that its verb representations are effectively grounded in the temporal dynamics of the input.
Improved unseen generalization.
In Figure 5 (a), we present the learning curve of our RCORE and the baseline [16] on Sth-com. Addressing the two key challenges in ZS-CAR, RCORE significantly improves unseen composition accuracy upon the baseline. While the baseline significantly increases FCP ratio from 10.8% at 5 epochs to 16.8% at 25 epochs, we observe our RCORE reports FCP ratio values of 10.8% at 5 epochs and 8.7% at 25 epochs, demonstrating stable mitigation of the co-occurrence bias. Notably, we also see a reduction in the seen-unseen accuracy gap.
Table 4: Ablation study on the Sth-com [16]. In every experiment, we use C2C [16] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). We report the performances using the unbiased open-world setting. The best H.M. numbers are highlighted.
(a) Effects of RCORE components. VOCAMix TORC Margin Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M. 63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.26 54.65 67.32 56.61 46.86 30.35 36.84 65.65 56.80 68.59 55.19 48.65 31.35 38.13 64.60 54.42 66.99 55.68 46.71 29.80 36.38 65.72 56.60 67.91 55.81 47.68 31.96 38.27 65.93 57.29 67.82 55.56 48.56 32.23 38.75
(b) Effects of TORC components. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.54 55.01 67.17 55.68 47.06 29.98 36.62 64.84 54.82 67.57 56.23 47.46 30.84 37.39 65.65 56.80 68.59 55.19 48.65 31.35 38.13
(c) Effects of VOCAMix. Method Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M. Baseline 63.60 54.36 67.72 56.10 46.31 30.08 36.47 CutMix [38] 63.87 52.40 68.08 57.01 46.63 29.00 35.76 Mixup [41] 62.79 52.42 67.84 56.34 45.82 29.37 35.79 VOCAMix w/o F.E. 64.26 54.35 67.61 56.19 46.82 30.07 36.62 VOCAMix 64.26 54.65 67.32 56.61 46.86 30.35 36.84
(d) Effects of Margin loss. Method Verb Object Composition FCP Seen Unseen Seen Unseen Seen Unseen H.M. Baseline (CE) 63.60 54.36 67.72 56.10 46.31 30.08 36.47 15.30 L-softmax [19] 63.39 52.87 67.56 55.96 46.59 29.11 35.84 10.93 Margin (Ours) 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
@@ 5.3 Comparisons with State-of-the-Art
Sth-com results.
As shown in Table 2, RCORE delivers consistent improvements across all metrics under both unbiased and biased evaluation on the Sth-com [16] dataset. Importantly, RCORE achieves a positive compositional gap $(\Delta_{\mathrm{CG}})$ on unseen compositions under the open-world unbiased setting, whereas all baselines exhibit negative gaps. This highlights the strong compositional generalization ability of RCORE. The superior verb representations learned by RCORE—validated in Section 5.2—translate directly into substantial gains on the ZS-CAR dataset.
EK100-com results.
In Table 3, RCORE consistently outperforms C2C [16] on both seen and unseen compositions under both unbiased and biased evaluation settings, despite the egocentric challenges. Notably, RCORE achieves a substantial improvement in unseen verb accuracy compared to C2C (52.07% vs. 49.71%), underscoring its stronger temporal modeling ability and reduced object-driven shortcuts.

### v2

@@ 4.3 Training
Component loss. We compute verb logits $\mathbf{S}_{V}\in\mathbb{R}^{|\mathbb{Y}^{V}|}$ and object logits $\mathbf{S}_{O}\in\mathbb{R}^{|\mathbb{Y}^{O}|}$ based on the cosine similarity scores. We apply standard cross-entropy losses for verb and object prediction: $L_{\text{com}}=L_{V}+L_{O}$.
Composition loss. Following prior work [li2024c2c], we obtain the composition logits $\mathbf{S}_{C}\in\mathbb{R}^{|\mathbb{Y}^{V}\times\mathbb{Y}^{O}|}$ by aggregating factorized logits: $\mathbf{S}_{C}=\mathbf{S}_{V}\boxplus\mathbf{S}_{O|V}+\mathbf{S}_{O}\boxplus\mathbf{S}_{V|O}$, where $\mathbf{S}_{O|V}$ and $\mathbf{S}_{V|O}$ represent the conditional logits and $\boxplus$ denotes broadcasted addition into the joint verb–object space. Then we employ the cross-entropy loss integrated with our batch-adaptive label space expansion (Section 4.1). Given $\mathbf{s}_{c}\in\mathbf{S}_{C}$, the final composition loss utilizing the dynamically constructed denominator $\mathbb{Y}_{\text{exp}}$ is formulated as: $L_{\text{comp}}=-\log(\exp(\hat{s}_{c})/\sum_{c=1}^{|\mathbb{Y}_{\text{exp}}|}\exp(s_{c}))$.
@@
Total loss. We train RCORE with the combined objective $L_{\text{total}}$ defined as $L_{\text{total}}=\alpha L_{\text{com}}+\beta L_{\text{comp}}+\gamma L_{\text{TORC}},+\delta L_{\text{CPR}},$ where $\alpha$, $\beta$, $\gamma$ and $\delta$ are hyperparameters.
@@ 5 Experimental Results
Building on our diagnosis, we evaluate whether RCORE reduces shortcut reliance and improves compositional generalization. We first describe the experimental setup (Section 5.1), then present diagnostic analyses (Section 5.2), followed by quantitative comparisons (Section 5.3), and finally ablation studies (Section 5.4).
@@ 5.1 Experimental setup
Limitations of conventional evaluation. Existing ZS-CAR works [li2024c2c, jiang2025dhd, jung2025crr] rely on the closed-world protocol, where inference is restricted to compositions that appear in the validation/test split. This protocol hides a model’s tendency to over-predict seen compositions and obscures co-occurrence-driven behaviors. More critically, test-set–tuned bias calibration selects a bias term using test labels, inflating unseen accuracy and undermining fair comparison.
@@
Our evaluation setup. To expose genuine generalization, we adopt an open-world, unbiased protocol by default: (i) inference spans the full space $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, and (ii) no test labels are used to tune biases. We also report closed-world H.M. and AUC for comparison with existing methods. We show open-world biased results using a realistic validation set-tuned bias in Section 0.E.4.
@@
Datasets. We evaluate on two ZS-CAR benchmarks: Sth-com [li2024c2c] and our curated EK100-com. Sth-com is derived from Something-Something V2 [goyal2017something] and Something-Else [mat2020sthelse], containing 79K videos with 161 verbs and 248 objects. Following prior work [yun2022time, laura2021temporal], we also evaluate on the Temporal subset consisting of verbs requiring temporal reasoning. We additionally introduce the EPIC-KITCHENS-100-composition (EK100-com) dataset, constructed by repurposing EK100 [damen2022ek100] following the Sth-com protocol. EK100-com includes 71K egocentric videos, 82 verbs, and 228 objects. Due to the long-tailed label distribution of EK100, it exhibits severe compositional sparsity. Its label coverage ratio is only $7.5\%$, which is considerably lower than the $12.8\%$ observed in Sth-com [li2024c2c]. Full details are in Section 0.B.
Figure 5: Analysis on the effects of RCORE on the Sth-com [li2024c2c] dataset. (a) RCORE suppresses the growth of FSP/FCP during training, unlike the baseline. (b) This reduces the seen–unseen gap and improves unseen composition accuracy. (c) On the temporal subset, RCORE shows a larger drop under temporal shuffling, indicating stronger temporal dependence over static cues. Best viewed with zoom and color.
Evaluation metrics. We report top-1 composition, verb and object accuracies. Verb and object accuracies are conditioned on the seen/unseen composition: verb@seen-comp and object@seen-comp are computed on samples with $\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}$, while verb@unseen-comp and object@unseen-comp are computed on samples with $\mathbf{y}^{C}\in\mathbb{Y}_{\text{unseen}}$. We report the corresponding harmonic mean (H.M.) between seen-comp and unseen-comp accuracies for composition, verb, and object. We also report the compositional gap $\Delta_{\mathrm{CG}}$ (2). For a detailed illustration of our evaluation setting, please refer Section 0.A.
@@
Baselines. We employ two encoder-based VLMs, CLIP-B/16 [radford2021clip] and InternVi-deo2-CLIP-B/14 [wang2024iv2], as backbones for ZS-CAR methods. We put the InternVid-eo2-1B[wang2024iv2] backbone results in the supplementary material. We primarily compare RCORE against C2C [li2024c2c], the current SOTA ZS-CAR method. To isolate the effect of conditional learning, we also evaluate the independent modeling baselines with CLIP [radford2021clip] and InternVideo2 [wang2024iv2] backbones which are denoted as ‘AIM’ and ‘LoRA’, named after their respective adapter tuning methods, in Section 5.3. AIM [yang2023aim] is the CLIP [radford2021clip]-based spatio-temporal adapter model. We additionally include closed-world results of a recent ZS-CAR approach [jung2025crr].
@@ 5.2 Analysis
We empirically show that RCORE reduces reliance on co-occurrence priors and mitigates object-driven shortcuts in verb learning, yielding more robust compositional behavior than the baseline [li2024c2c]. Here, we present results with the CLIP [radford2021clip] backbone, and we put InternVideo2 [wang2024iv2] results in the supplementary.
C2C (Baseline)      RCORE (Ours)
Figure 6: RCORE mitigates object-driven shortcuts in verb learning. We visualize confusion matrices for six representative verbs on unseen compositions of the Sth-com [li2024c2c] test set. All values are normalized frequencies across the six verb classes.
@@
RCORE less relies on co-occurrence prior. In Figure 5 (a), we present the False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP) curves of RCORE and the baseline [li2024c2c] on Sth-com. While the baseline’s FSP and FCP increase substantially during training (53% $\rightarrow$ 63% and 26% $\rightarrow$ 29%), those of RCORE remain consistently lower and even decrease (47% $\rightarrow$ 44% and 23% $\rightarrow$ 23%), indicating markedly less dependence on co-occurrence statistics.
Improved unseen generalization. In Figure 5 (b), we present the learning curve of our RCORE and the baseline [li2024c2c] on Sth-com. RCORE significantly improves unseen composition accuracy upon the baseline (34% vs. 30%), and reduces the seen–unseen accuracy gap (9 points vs. 17 points).
@@
The learned verb representations are indeed effective. Inspired by prior work [yun2022time], we evaluate models on the Temporal subset of Sth-com [li2024c2c], as shown in Figure 5 (c). Specifically, to measure a model’s dependence on static cues, we compare its performance when using temporally shuffled verb features $\mathbf{f}^{V}_{\text{shuffled}}$ versus the original features $\mathbf{f}^{V}$. Across both seen and unseen compositions, RCORE exhibits a substantially larger performance gap between original and shuffled features. This demonstrates that its verb representations are effectively grounded in temporal dynamics.
In Figure 6, we visualize verb confusion matrices on unseen compositions of the Sth-com test set. RCORE discriminates pairs of opposite verbs, while C2C [li2024c2c] frequently confuses them, e.g. misclassifying ‘unfolding’ as ‘folding’. These results collectively demonstrate that C2C fails to overcome the challenge of object-driven shortcuts in verb learning, whereas RCORE effectively mitigates it.
Table 2: Sth-com [li2024c2c] results. We show the top-1 verb, object, and composition classification accuracy (%). We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Backbone Method Verb Object Composition
@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
CLIP AIM [yang2023aim] 54.99 39.19 45.76 67.19 53.43 59.53 40.24 (+3.29) 18.50 (-2.44) 25.35
C2C [li2024c2c] 63.60 54.36 58.62 67.72 56.10 61.36 46.31 (+3.24) 30.08 (-0.42) 36.47
RCORE 65.73 59.00 62.18 64.79 56.34 60.27 44.99 (+2.40) 33.90 (+0.66) 38.67
InternVideo2 LoRA [hu2022lora] 48.83 36.50 41.78 64.79 58.61 61.55 32.83 (+1.19) 19.71 (-1.68) 24.63
C2C [li2024c2c] 70.35 63.29 66.63 70.03 63.44 66.58 51.50 (+2.24) 39.53 (-0.63) 44.73
RCORE 71.65 66.65 69.06 67.96 64.56 66.22 50.20 (+1.51) 43.98 (+0.95) 46.88
@@ 5.3 Comparisons with State-of-the-Art
Sth-com. As shown in Table 2, RCORE achieves clear gains on Sth-com [li2024c2c] under a rigorous open-world setting. Compared to the C2C [li2024c2c] baseline, RCORE improves the verb@unseen-comp by 4.6 points and the unseen composition accuracy by 3.8 points with the CLIP [radford2021clip] backbone, and by 3.4 points and 4.5 points with InternVideo2 (IV2) [wang2024iv2], respectively. These gains are consistent with stronger verb recognition observed in our analysis (Section 5.2). Notably, RCORE yields a positive compositional gap ($\Delta_{\text{CG}}$) on unseen compositions, whereas all baselines remain negative, indicating improved compositional behavior on unseen pairs. This improvement coincides with the highest composition H.M. under the open-world protocol. In Table 3, RCORE also records the highest H.M. and AUC under the closed-world, test-set-tuned protocol used in prior work [jung2025crr].
Table 3: Closed-world results on Sth-com[li2024c2c].
Method Closed-world
H.M. AUC
AIM [yang2023aim] 34.44 16.99
C2C [li2024c2c] 42.75 23.77
Jung et al. [jung2025crr] 42.60 23.50
RCORE 43.21 25.27
EK100-com. Despite the strong co-occurrence prior in the EK100-com training compositions (see Section 5.1), RCORE achieves higher unseen composition accuracy and the best H.M. As shown in Table 4, compared to C2C [li2024c2c], RCORE substantially improves unseen composition accuracy—by 6.9 points with the CLIP [radford2021clip] backbone and 7.7 points with the IV2 [wang2024iv2] backbone—despite a decrease in seen composition accuracy. This pattern is consistent with reduced reliance on seen co-occurrence statistics and improved recognition of unseen compositions. Moreover, RCORE yields a positive compositional gap ($\Delta_{\text{CG}}$) on unseen compositions (+1.32 with CLIP and +3.68 with IV2), which is consistent with improved compositional behavior under highly sparse supervision.
Table 4: EK100-com results. We show the top-1 verb, object, and composition classification accuracy (%). We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Backbone Method Verb Object Composition
@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
CLIP AIM [yang2023aim] 59.97 39.99 47.98 56.23 46.76 51.06 38.58 (+4.86) 14.98 (-3.71) 21.58
C2C [li2024c2c] 65.52 49.57 56.44 56.54 46.99 51.32 42.70 (+5.65) 21.56 (-1.73) 28.65
RCORE 66.19 54.31 59.66 54.35 49.88 52.02 39.70 (+3.73) 28.41 (+1.32) 33.12
InternVideo2 LoRA [hu2022lora] 59.73 32.99 42.51 51.08 39.34 44.45 36.43 (+5.92) 5.63 (-7.35) 9.75
C2C [li2024c2c] 68.95 57.01 62.41 60.54 52.31 56.12 47.30 (+5.55) 30.33 (+0.51) 36.96
RCORE 69.00 60.98 64.74 56.59 56.40 56.49 42.32 (+3.27) 38.07 (+3.68) 40.08